# SpotV2Net: Multivariate Intraday Spot Volatility Forecasting via Vol-of-Vol-Informed Graph Attention Networks

### v2

@@ 7.2 Single-step forecast
We compare the forecasting accuracy of SpotV2Net with that of three alternative models, detailed in App. A: the panel HAR-Spot model, the XGB model and LSTM model. We also compare the forecasting accuracy of SpotV2Net with that of a modified version of the latter, which does not accept edge features as input but only node features. We refer to it as SpotV2Net with no edge features, shortened to SpotV2Net-NE. Tab. 2 shows aggregate values of Mean Squared Error (MSE), that is,
@@
and QLIKE, that is,
$\frac{1}{T_{\text{test}}}\sum_{j=1}^{T_{\text{test}}}\frac{1}{N}\sum_{i=1}^{N}\left[\frac{\textrm{exp}\{\widehat{V}_{i,j}\}}{\textrm{exp}\{\widehat{V}^{*}_{i,j}\}}-\left(\widehat{V}_{i,j}-\widehat{V}^{*}_{i,j}\right)-1\right],$ (3)
@@
Aggregate performance measures in Tab. 2 suggest that SpotV2Net outperforms the HAR-Spot model and the other ML models considered, namely XGB and LSTM, both during the validation and test periods. Further, while the LSTM and XGB models represent an improvement in forecasting accuracy compared to the HAR-Spot model in the validation period, the latter shows a better performance, compared to LSTM and XGB, in the test period. SpotV2Net-NE is less accurate than SpotV2Net in both the validation and the test phases. Moreover, it is worth noting that SpotV2Net-NE still outperforms HAR-Spot and the other ML methods in the validation period. Instead, in the test period, SpotV2Net-NE performs worse than HAR-Spot, but still better than LSTM and XGB. These findings underline the crucial role of edge features (that is, the values of the spot co-volatility of volatility matrix) in capturing spillover effects and improving forecasting accuracy. Note that the improved performance by all models in the validation period is due to the fact that the validation employs data for fine-tuning model hyperparameters.
Tab. 2 also reports the results of the Model Confidence Set (MCS) (Hansen et al.,, 2011), which is applied column-wise to identify a subset of models that demonstrate superior performance at the 5% significance level. The MCS results show that SpotV2Net is the only model that remains in the confidence set at the $5$% significance level and validate its superior performance against the alternative forecasting models employed in our study.
@@
Table 3: Comparison of single-step forecasting accuracy using the Diebold-Mariano (DM) test with MSE (top) and QLIKE (bottom) losses in the validation period. Positive numbers indicate the column model outperforms the row model. Superscript * denotes the significance levels of 5%
@@
Table 4: Comparison of single-step forecasting accuracy using the Diebold-Mariano (DM) test with MSE (top) and QLIKE (bottom) losses in the test period. Positive numbers indicate the column model outperforms the row model. Superscript * denotes the significance levels of 5%
@@ 7.3 Model Interpretation
For this purpose, we employ GNNExplainer (Ying et al.,, 2019). GNNExplainer is a model-agnostic approach that provides interpretable explanations for predictions of GNN-based models. Specifically, GNNExplainer allows the identification of the subgraph [^10] A subgraph is defined as a graph consisting of a subset of the nodes and edges of the original graph. deemed the most influential for the model’s prediction of a specific node.
@@
For the interpretation of the single-step forecasts produced by SpotV2Net for the spot volatilities of the DIJA constituents, we implement GNNExplainer for each of the $N=30$ nodes (i.e., asset spot volatilities) of the input graph, with the aim of extrapolating, for each node, a subgraph composed of the $N^{*}=5$ nodes representing the most influential nodes for predicting the node’s spot volatility. The implementation is repeated at each time step [^11] For the implementation, we use the Python package available at the link GNNExplainer. Details about the optimization algorithm are provided in Section 4 of Ying et al., (2019)..
@@ 7.4 Multi-step forecast
Accurate multi-step forecasting can be advantageous for high-frequency traders, who wish not only to predict volatilities at the immediate next time step but also to extrapolate entire trajectories, enabling them to plan their operations with a broader, daily, temporal perspective. Accordingly, in this Section, we extend the scope of application of the SpotV2Net model to obtain multiple-step-ahead forecasts. Specifically, we adapt the model’s architecture to predict the entire daily volatility trajectory on the 30-minute grid, which corresponds to 14 future observations for a 6.5-hour long trading day [^12] In principle, it is possible to adapt SpotV2Net’s architecture to produce multi-step forecasts at any point in time. For simplicity, we restrict our study to multi-step forecasts obtained at the end of a trading day for the entire next trading day.. Note that multi-step forecasts are all conditional on the information available at the end of the previous trading day (i.e., node and edge features of the graph at the end of the previous trading day). In other words, letting $\tau_{0}$ denote the end of the previous day, the multi-step version of SpotV2Net does not use as input the predictions at times $\tau_{1},...,\tau_{h-1}$ to obtain the prediction at time $\tau_{h}$, $1<h\leq 14$. Instead, the model produces, at $\tau_{0}$, a unique joint forecast of the values of the spot volatility function on the time grid $\tau_{1},...,\tau_{14}$, with mesh size equal to $30$ minutes [^13] The ability to produce a unique multi-step joint forecast is characteristic to deep learning models, like SpotV2Net and the LSTM model. Such an ability is not shared by the HAR-Spot model and the XGB model. Indeed, to obtain multi-step forecasts with these two models, one has to resort to a recursive approach, where the one-step ahead forecast for time $\tau_{j}$ is employed as an input to obtain the one-step ahead forecast for time $\tau_{j+1}$..
We evaluate the multi-step SpotV2Net against the same alternative models considered in Sec. 7.2. For the evaluation of multi-step forecasts, MSE is modified as
$\frac{1}{D_{\text{test}}}\sum_{j=1}^{D_{\text{test}}}\frac{1}{N}\sum_{i=1}^{N}\frac{1}{H}\sum_{h=0}^{H}\left(\widehat{V}_{i,j+h}-\widehat{V}^{*}_{i,j+h}\right)^{2}$ (5)
and, similarly, QLIKE is modified as
$\frac{1}{D_{\text{test}}}\sum_{j=1}^{D_{\text{test}}}\frac{1}{N}\sum_{i=1}^{N}\frac{1}{H}\sum_{h=0}^{H}\left[\frac{\textrm{exp}\{\widehat{V}_{i,j+h}\}}{\textrm{exp}\{\widehat{V}^{*}_{i,j+h}\}}-\left(\widehat{V}_{i,j+h}-\widehat{V}^{*}_{i,j+h}\right)-1\right]$ (6)
@@
where $H$ denotes number of forecast steps and $D_{test}$ indicates the number of days in the test set. When the validation period is considered, $D_{\text{test}}$ is replaced by $D_{\text{val}}$, which indicates the number of days in the validation period. Note that we set $H=14$ in our study.
@@
Tab. 5 summarizes the multi-step forecasting performance, in the validation and test phases, of the models employed in our study. In the validation period, the ranking is the same as for single-step forecasts, both in terms of aggregate values of MSE and QLIKE. Instead, in the test period, the HAR-Spot model has a worse ranking than in the single-step exercise, as it outperforms only the XGB model. This finding is in line with the results by Christensen et al., (2022), which suggest that the relative forecasting accuracy of HAR-type models may deteriorate at longer forecasting horizons, compared to that of ML models. Tab. 5 reports also the results of the application of the MCS at the 5% significance levels. The outcome is analogous to the one obtained in the single-step case and supports the superior forecasting performance of SpotV2Net also in the multi-step exercise.
@@
Moreover, the DM test statistics are reported in Tabs. 6 and 7. The results suggest that the gain in forecasting accuracy provided by SpotV2Net is statistically significant at the 5% level also for multi-step forecasts.
@@
Table 6: Comparison of multi-step forecasting accuracy using the Diebold-Mariano (DM) test with MSE (top) and QLIKE (bottom) losses on the validation set. Positive numbers indicate the column model outperforms the row model. Superscript * denotes the significance levels of 5%.
@@
Table 7: Comparison of multi-step forecasting accuracy using the Diebold-Mariano (DM) test with MSE (top) and QLIKE (bottom) losses on the validation set. Positive numbers indicate the column model outperforms the row model. Superscript * denotes the significance levels of 5%.
@@ 8 Conclusions
This paper contributes to research on volatility forecasting by introducing SpotV2Net, a model that employs a neural network with a GAT architecture to produce intraday forecasts of multivariate spot volatilities. Specifically, SpotV2Net leverages the interconnected nature of financial markets by modeling each asset as a node in a graph structure, where node and edge features represent, respectively, univariate and multivariate spot volatilities and volatilities of volatilities. Further, the attention mechanism embedded in the GAT architecture allows SpotV2Net to optimally weigh the influence of specific subgraphs when producing spot volatility forecasts for a specific node.
@@ A.3 Long Short-term Memory (LSTM) Neural Network
where $\mathbf{x}_{\tau_{b}}$ is the input vector, $\mathbf{f}_{\tau_{b}}$ is the forget gate’s activation vector, $\mathbf{i}_{\tau_{b}}$ is the update gate’s activation vector, $\mathbf{o}_{\tau_{b}}$ is the output gate’s activation vector, $\tilde{\mathbf{c}}_{\tau_{b}}$ is the cell input activation vector, $\mathbf{c}_{\tau_{b}}$ is the cell state vector and $\mathbf{h}_{\tau_{b}}$ is the hidden state vector, i.e. the output vector of the LSTM unit. Further, $\circ$ is the Hadamard product function, $\sigma_{g}$ is the sigmoid function and $\sigma_{c}$ and $\sigma_{h}$ are hyperbolic tangent functions. Finally, $\mathbf{W}_{(\cdot)}$ and $\mathbf{b}_{(\cdot)}$ refer to weight matrices and bias vectors that need to be estimated by training the model, while $\mathbf{U}_{(\cdot)}$ refers to the weight matrices that are applied to the hidden state vector from the previous time step $\mathbf{h}_{\tau_{b-1}}$.

### v3

@@ 7.2 Single-step forecast
We compare the forecasting accuracy of SpotV2Net with that of three alternative models, detailed in App. A: the panel HAR-Spot model, the XGB model and LSTM model. We also compare the forecasting accuracy of SpotV2Net with that of a modified version of the latter, which does not accept edge features as input but only node features. We refer to it as SpotV2Net with no edge features, shortened to SpotV2Net-NE. Tab. 2 shows aggregate values of the Mean Squared Error (MSE), that is,
@@
and the QLIKE, that is,
${\color[rgb]{0,0,0}\frac{1}{T_{\text{test}}}\sum_{j=1}^{T_{\text{test}}}\frac{1}{N}\sum_{i=1}^{N}\left[\frac{\widehat{V}_{i,j}}{\widehat{V}^{*}_{i,j}}-\log\frac{\widehat{V}_{i,j}}{\widehat{V}^{*}_{i,j}}-1\right]},$ (5)
@@
Aggregate performance measures in Tab. 2 suggest that SpotV2Net outperforms the HAR-Spot model and the other ML models considered, namely XGB and LSTM, both during the validation and test periods. Besides, while the LSTM and XGB models represent an improvement in forecasting accuracy compared to the HAR-Spot model in the validation period, the latter shows a better performance, compared to LSTM and XGB, in the test period. SpotV2Net-NE is less accurate than SpotV2Net in both the validation and the test phases. Moreover, it is worth noting that SpotV2Net-NE still outperforms HAR-Spot and the other ML methods in the validation period. Instead, in the test period, SpotV2Net-NE performs worse than HAR-Spot, but still better than LSTM and XGB. These findings underline the crucial role of edge features (that is, the values of the spot co-volatility of volatility matrix) in capturing spillover effects and improving forecasting accuracy. Note that the improved performance by all models in the validation period is due to the fact that the validation employs data for fine-tuning model hyperparameters.
Tab. 2 also reports the results of the Model Confidence Set (MCS) (Hansen et al.,, 2011), which is applied column-wise to identify a subset of models that demonstrate superior performance with a 95% level of confidence. We implemented the MCS with 5000 bootstrap replications, using the range statistic to compute the p-values. The results show that the subset of superior models contains only SpotV2Net, indicating that the latter yields superior predictive performance, compared to the other models considered, with a level of confidence of 95%.
@@
Table 3: Comparison of single-step forecasting accuracy using the Diebold-Mariano (DM) test with MSE (top) and QLIKE (bottom) losses in the validation period. Positive numbers indicate the column model outperforms the row model. Superscript * denotes the significance levels of 5%.
@@
Table 4: Comparison of single-step forecasting accuracy using the Diebold-Mariano (DM) test with MSE (top) and QLIKE (bottom) losses in the test period. Positive numbers indicate the column model outperforms the row model. Superscript * denotes the significance levels of 5%.
@@ 7.3 Model Interpretation
For this purpose, in this Section we employ GNNExplainer (Ying et al.,, 2019), a model-agnostic approach that provides interpretable explanations for predictions of GNN-based models. Specifically, GNNExplainer allows the identification of the subgraph [^10] A subgraph is defined as a graph consisting of a subset of the nodes and edges of the original graph. deemed the most influential for the model’s prediction of a specific node.
@@
For the interpretation of the single-step forecasts produced by SpotV2Net for the spot volatilities of the DIJA constituents, we implement GNNExplainer for each of the $N=30$ nodes (i.e., asset spot volatilities) of the input graph, with the aim of extrapolating, for each node, a subgraph composed of the $N^{*}=5$ nodes representing the most influential nodes for predicting the node’s spot volatility. Note that we use GNNEXPLAINER to identify the most influential subgraph, not to isolate the most influential individual features of specific nodes. This approach provides a global perspective on the relevance of a particular subgraph based on the entire information set embedded in its nodes and edges. This way, GNNExplainer complements traditional econometric methods, which typically emphasize the statistical significance of the contribution of individual features. The implementation is repeated at each time step [^11] For the implementation, we use the Python package available at the link GNNExplainer. Details about the optimization algorithm are provided in Section 4 of Ying et al., (2019)..
@@ 7.4 Multi-step forecast
Accurate multi-step forecasting can be advantageous for high-frequency traders, who wish not only to predict volatilities at the immediate next time step but also to extrapolate entire trajectories, enabling them to plan their operations with a broader, daily, temporal perspective. Accordingly, in this Section, we extend the scope of application of the SpotV2Net model to obtain multiple-step-ahead forecasts. Specifically, we adapt the model’s architecture to predict the entire daily volatility function on the 30-minute grid, which corresponds to 14 future observations for a 6.5-hour long trading day  [^12] In principle, it is possible to adapt SpotV2Net’s architecture to produce multi-step forecasts at any point in time. For simplicity, we restrict our study to multi-step forecasts obtained at the end of a trading day for the entire next trading day.. Note that multi-step forecasts are all conditional on the information available at the end of the previous trading day (i.e., node and edge features of the graph at the end of the previous trading day). In other words, letting $\tau_{0}$ denote the end of the previous day, the multi-step version of SpotV2Net does not use as input the predictions at times $\tau_{1},...,\tau_{h-1}$ to obtain the prediction at time $\tau_{h}$, $1<h\leq 14$. Instead, the model produces, at $\tau_{0}$, a unique joint forecast of the values of the spot volatility function on the time grid $\tau_{1},...,\tau_{14}$, with mesh size equal to $30$ minutes. We refer to this type of multi-step forecast as functional forecast. We evaluate the performance of SpotV2Net in producing such type of forecast against the same alternative models considered in Sec. 7.2.131313The ability to produce a functional forecast is characteristic of deep learning models, like SpotV2Net and the LSTM model. However, such an ability is not shared by the HAR-Spot model and the XGB model. Indeed, to obtain multi-step forecasts with these models, one has to resort to a recursive approach, where the one-step ahead forecast for time $\tau_{j}$ is employed as an input to obtain the one-step ahead forecast for time $\tau_{j+1}$..
To evaluate functional forecasts accuracy, we use $global$ loss functions. Specifically, we employ the average MSE of the multi-step prediction, that is,
$\frac{1}{D_{\text{test}}}\sum_{j=1}^{D_{\text{test}}}\frac{1}{N}\sum_{i=1}^{N}\frac{1}{H}\sum_{h=1}^{H}\left(\widehat{V}_{i,j+h}-\widehat{V}^{*}_{i,j+h}\right)^{2}$ (7)
and the average QLIKE, that is,
${\color[rgb]{0,0,0}\frac{1}{D_{\text{test}}}\sum_{j=1}^{D_{\text{test}}}\frac{1}{N}\sum_{i=1}^{N}\frac{1}{H}\sum_{h=1}^{H}\left[\frac{\widehat{V}_{i,j+h}}{\widehat{V}^{*}_{i,j+h}}-\log\frac{\widehat{V}_{i,j+h}}{\widehat{V}^{*}_{i,j+h}}-1\right],}$ (8)
@@
where $H=14$ denotes the number of intraday 30-minute forecast steps and $D_{test}$ indicates the number of days in the test set. When the validation period is considered, $D_{\text{test}}$ is replaced by $D_{\text{val}}$, which indicates the number of days in the validation period.
@@
Tab. 5 summarizes the multi-step forecasting performance of the models employed in our study, in the validation and test phases. In the validation period, the ranking is the same as for single-step forecasts, both in terms of aggregate values of average MSE and QLIKE. Instead, in the test period, the HAR-Spot model has a worse ranking than in the single-step exercise, as it outperforms only the XGB model. This finding is in line with the results by Christensen et al., (2023), which suggest that the relative forecasting accuracy of HAR-type models may deteriorate at longer forecasting horizons, compared to that of ML models. Tab. 5 reports also the results of the application of the MCS at the 95% confidence level. The outcome is analogous to the one obtained in the single-step case and supports the superior forecasting performance of SpotV2Net also in the multi-step exercise.
@@
Moreover, the DM test statistics are reported in Tabs. 6 and 7. The results suggest that the gain in forecasting accuracy provided by SpotV2Net is statistically significant at the 5% level also for multi-step forecasts. Finally, to illustrate an example of the accuracy of the multi-step functional predictions produced by SpotV2Net, Fig. 5 shows a comparison of the latter with the actual reconstructed spot volatility path of four companies for two days, one in the validation set and one in the test set.
@@
Table 5: Aggregate values of average MSE and QLIKE in Eqns. (7) and (8) for the SpotV2Net model and the alternative models during the validation and test phases. An asterisk (*) indicates models that are included in the MCS at the 5% significance level.
@@
Table 6: Comparison of multi-step forecasting accuracy using the Diebold-Mariano (DM) test with average MSE (top) and QLIKE (bottom) losses on the validation set. Positive numbers indicate the column model outperforms the row model. Superscript * denotes the significance levels of 5%.
@@
Table 7: Comparison of multi-step forecasting accuracy using the Diebold-Mariano (DM) test with average MSE (top) and QLIKE (bottom) losses on the validation set. Positive numbers indicate the column model outperforms the row model. Superscript * denotes the significance levels of 5%.
Figure 5: Predicted (dotted blue) and actual (reconstructed, dashed grey) spot volatility path for AXP, DIS, HON, and MSFT. The first row refers to September 19, 2022, comprised in the validation period, and the second row to February 1, 2023, comprised in the test period.
@@ 8 Conclusions
This paper contributes to research on volatility forecasting by introducing SpotV2Net, a model that employs a neural network with a GAT architecture to produce intraday forecasts of multivariate spot volatilities. Specifically, SpotV2Net leverages the interconnected nature of financial markets by modeling each asset as a node in a graph structure, where node and edge features represent, respectively, univariate and multivariate spot volatilities and volatilities of volatilities. Furthermore, the attention mechanism embedded in the GAT architecture allows SpotV2Net to optimally weigh the influence of specific subgraphs when producing spot volatility forecasts for a specific node.
@@ A.3 Long Short-term Memory (LSTM) Neural Network
where $\mathbf{x}_{\tau_{b}}$ is the input vector, $\mathbf{f}_{\tau_{b}}$ is the forget gate’s activation vector, $\mathbf{i}_{\tau_{b}}$ is the update gate’s activation vector, $\mathbf{o}_{\tau_{b}}$ is the output gate’s activation vector, $\tilde{\mathbf{c}}_{\tau_{b}}$ is the cell input activation vector, $\mathbf{c}_{\tau_{b}}$ is the cell state vector and $\mathbf{h}_{\tau_{b}}$ is the hidden state vector, i.e. the output vector of the LSTM unit. Besides, $\circ$ is the Hadamard product function, $\sigma_{g}$ is the sigmoid function and $\sigma_{c}$ and $\sigma_{h}$ are hyperbolic tangent functions. Finally, $\mathbf{W}_{(\cdot)}$ and $\mathbf{b}_{(\cdot)}$ refer to weight matrices and bias vectors that need to be estimated by training the model, while $\mathbf{U}_{(\cdot)}$ refers to the weight matrices that are applied to the hidden state vector from the previous time step $\mathbf{h}_{\tau_{b-1}}$.